OpenStudy (anonymous):

factor this equation. x squared -9x-36=0. show work

jimthompson5910 (jim_thompson5910):

x^2 - 9x - 36 = 0 x^2 - 12x + 3x - 36 = 0 (x^2 - 12x) + (3x - 36) = 0 x(x - 12) + 3(x - 12) = 0 (x + 3)(x - 12) = 0 x + 3 = 0 or x - 12 = 0 x = -3 or x = 12

jimthompson5910 (jim_thompson5910):

Take note that the two numbers -12 and 3 both multiply to -36 and add to -9.